A trapezoid has an area of 117 square feet and a height of 9 feet. One base is 4 feet shorter than the other base. How long is each base? Write an equation and solve it.
Answer by stanbon(75887)

A trapezoid has an area of 117 square feet and a height of 9 feet. One base is 4 feet shorter than the other base. How long is each base? Write an equation and solve it.
---------------------------
Formula: Area = (h/2)(b1 + b2)
117 = (9/2)(x + (x-4))
(2/9)117 = 2x-4
26 = 2x - 4
2x = 30
x = 15 ft. (one of the bases)
x-4 = 11 ft. (the other base)
=============================
Cheers,
Stan H.